OpenStudy (anonymous):

If f(x) = 5x + 4 and g(x) = 6x – 3, what is f(g(x))?

OpenStudy (anonymous):

I'm thinking 30x+21

OpenStudy (anonymous):

f(x) = 5x + 4 and g(x) = 6x – 3 f(g(x))=5(6x – 3)+ 4 f(g(x))=30x -15+ 4 f(g(x))=30x -11

OpenStudy (anonymous):

plug g into f so f(g(x))=5(6x-3)+4

OpenStudy (anonymous):

close good job though